A random sample of 30 pair of observations from a normal population gives a correlation coefficient of 0.32 . Is it likely that the variable in the coefficient are correlated?

1
Expert's answer
2021-12-13T14:23:25-0500

"H_0:r=0" variables in the coefficient are not correlated

"H_a:r\neq0" variables in the coefficient are correlated


"t=\frac{r\sqrt{n-2}}{1-r^2}=\frac{0.32\sqrt{30-2}}{1-0.32^2}=1.886"


"df=n-2=30-2=28"

for "\alpha=0.05"

critical value:

"t_{crit}=2.048"


Since "|t|<t_{crit}" we accept null hypothesis. Variables in the coefficient are not correlated.